-#1 points SerCP10 13.P046 A bat can detect small objects, such as an insect, whose size is approximately equal to 7.99 x 10* Hz, and if the speed of sound in air is 343 m/s, what is the smalest insect a bat can detect? My Notes Ask Your Teack one wavelength of the sound the bat makes. If bats emit a chirp at a frequency of Need Help? Read ot points | Previous Answers SerCP10 13 P061 My Notes Ask Your A pilano string having a mass per unit length equal to 4.90 x 10-3 kg/m is under a tension of 1,250 N. Find the speed with which a wave travels on this string 494.87 Your response is within 10% of the correct value. This least four-digit accuracy to minimize roundoff error, mys may be due to undo em oryou could have stake in your ca la Cany out al inter edate re stoat Need Holp? Pek
Please help with any or all problems I can’t seem to figure them out thanks
Answer #1

1) Smallest insect length = 1 wavelength

wavelength = velocity/ frequency
= 343 m/s/ 7.99x10^4Hz

length = 4.29E-3 m = 4.29 mm

Smallest insect bat can detect = 4.29 mm

2) Speed = sqrt (tension/mass per unit length) = sqrt(1250/4.9*10^-3) = 505.07 m/s